Day shift summary Nov. 14, 1999

Taku_Yamanaka (taku@fnal.gov)
Sun, 14 Nov 1999 15:29:36 -0600

Shift Summary for DAY shift, Nov 14, 1999
========================================================================
On Shift: Taku & Carlos
========================================================================
Hours of beam available at all intensities: 5.7h
Hours of beam available at nominal intensity: 4.5h
Hours of E799 data (special runs not included): 4.7h
Hours of special runs (laser scans, muon runs...): 0.9h
========================================================================
Accounting of lost time: KFREND/fighting with crate 53 : 40min
========================================================================
Runs ended this shift: Run 14962 - 234K 3pi0 Dal
Run 14963 - Laser scan
Run 14964 - Muon run
========================================================================
Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: None
========================================================================
Overview of shift: E799 physics data.
========================================================================
Description of Shift:
Problems
csiprobs ch 1096 gain has drifted down by 6.3%.
BA BA ADC counts peaks below at below 0.
Muon Mu2, Mu3x, Mu3y timings are -1.53 - -1.59 ns.
scia SA4 ADC mean is ~10, instead of ~30.
dc 3x has sigma = 0.205 > 0.2.

7:01 Taku arrived.
8:43 Linac problem
9:28 Stopped Run 14962 BEAM_799.
9:36 Ramped down AN4 (AN799POSTONEG)
9:48 Start Run 14963 Laser scan.
10:08 Switched NM2SR from NEG to POS.
10:30 Stop Run 14963.
10:34 Update CsI constant (CSQ614963).
11:34 Beam returns.
Start Run 14964 Muon Run.
11:48 Stop Run 14964. #Trig1 > 100k.
11:50 Turned on AN4 (NEG).
11:58 Start Run 14965 BEAM_799, but crate 53 keeps on hanging up.
Lots of reseating grey cable, KFRENDing, recycling
powers, ...
12:20 Stop Run 14965. Don't use this run.
Update T0 cosnt (DCT014961) made by Ashkan.
12:30 Start Run 14966 BEAM_799
KUMQUAT hung. Reset it.
12:38 Finally, smooth data taking.
~13:45 Intensity has come back to 8E12.

========================================================================
Follow up Notes on Previous Shift: none
========================================================================
Items to Follow up on Next Shift: none
========================================================================